OpenStudy (cicilybailey14):

Write an equation in point-slope form for the line through the given point with the given slope. (4, 7), m = 3/2 y - 7 = 3/2(x - 4) y = 3/2x -3 y - 4 = 3/2(x - 7) y = 3/2x + 3

OpenStudy (godlovesme):

Y-y1=m(x-x1) M is slope Y1 is 7 X1 is 4 Plug them into the eqn

OpenStudy (anonymous):

if line passes through point (x1,y1) and with slope m yhen equation of line is (y-y1)=m(x-x1) substitue the given values and you will find ans. y-7=3/2(x-4)